OpenStudy (anonymous):

PLLLLEEEEZZ HELP!! ASAP! <3 Part 1 (4 points)] Show all work in simplifying (3x2+3x-36)/(x 2-x-6) Be sure to list restrictions. [Part 2 (2 points)] Use complete sentences to explain your process

OpenStudy (anonymous):

\[(3x^2+3x-36)/(x^2-x-6)\] \[=(3x-9)(x+4)/(x-3)(x+2)\] \[=3(x-3)(x+4)/(x-3)(x+2)\] \[=3(x+4)/(x+2), x \neq -2\]

OpenStudy (anonymous):

Can you please explain how you did this?? in complete sentances!? @2InfinityAndBeyond

OpenStudy (anonymous):

... anybody!

OpenStudy (anonymous):

1. factor the numerator and denominator 2. cancel the common factors from the numerator and denominator